The inductive reactance of a 20.0-mH inductor at a certain frequency is 120 Ω. What is the frequency in Hz? A) 6000 Hz B) 120 Hz C) 478 Hz D) 955 Hz E) 167 Hz
Answer #1

apply inductive reactance XL = 2pifL

so f = 120/(2*3.14*0.02))

f = 955 Hz

option D it is
Answer #2

955 hz
Answer #3

reactance=2?fL

120=2*3.14*f*20.0*10-3

f=955 Hz

Hence D. 955 Hz
Answer #5

lmsin(wt -90。)

L=20*10^-3H

XL=120 ohm

omega =2*pi*f

f=XL/(2pi*L)

f = 120/(2*3.14*0.02))

f = 955 Hz (option D)